OpenStudy (angela210793):

Some help?? plsss... In a human DNA molecule 30% is adenine. Find the % of guanine, cytosine and thymine...

OpenStudy (anonymous):

% of adenine = % of thymine=30% rest(100-60)% is contributed by cytosine and guanine % of cytosine = % of Guanine % of cytosine = 20% % of guanine = 20% ref : Chargaff's rule I think this is the way it is.............

OpenStudy (angela210793):

thnx :)

OpenStudy (anonymous):

A=T SO 30+30=60 NOW REST IS C+G WHICH IS 40 ,SO,C=G SO40/2=20 EACH